### v1

# On a problem of Nathanson on non-minimal additive complements

Abstract.
Let $C$ and $W$ be two sets of integers. If $C+W=\mathbb{Z}$, then $C$ is called an additive complement to $W$. We further call $C$ a minimal additive complement to $W$ if no proper subset of $C$ is an additive complement to $W$. Answering a problem of Nathanson in part, we give sufficient conditions of $W$ which has no minimal additive complements. Our result also extends the prior result of Chen and Yang.
Key words and phrases:
integer sets; minimal additive complement
2010 Mathematics Subject Classification:
11B13
ORCID: $0$-0002-0439-6079
†† ORCID: $0$-0001-7016-309X  &\&&  Corresponding author
1. Introduction
Let $\mathbb{Z}$ be the set of all integers. For $C,W\subseteq\mathbb{Z}$ and $c\in\mathbb{Z}$, define
$C+W:=\big{\{}c+w:c\in C,~{}w\in W\big{\}}\quad\text{and}\quad cW:=\big{\{}cw:c\in C\big{\}}.$
For the set $W=\{w\}$ consisted of a single element, we just write $C+\{w\}$ as $C+w$ for simplicity. If $C+W=\mathbb{Z}$, then $C$ is called an additive complement to $W$. Furthermore, if any proper subset of $C$ is not an additive complement to $W$, then $C$ is called a minimal additive complement to $W$. Sometimes, for simplicity, a minimal additive complement is just called a minimal complement throughout this article.
In 2011, Nathanson [9, Theorem 8] proved the following interesting theorem.
Nathanson’s theorem.
Let $W$ be a nonempty, finite set of integers. In $\mathbb{Z}$, every complement to $W$ contains a minimal complement to $W$.
Nathanson [9, Problem 11] also posed the following open problem.
Problem 1.
Let $W$ be an infinite set of integers. Does there exist a minimal complement to $W$? Does there exist a complement to $W$ that does not contain a minimal complement?
In 2012, Chen and Yang [5, Theorem 1] considered Problem 1. For the situation $\inf W=-\infty$ and $\sup W=+\infty$, they made the following progress.
Theorem A (Chen–Yang).
Let $W$ be a set of integers with $\inf W=-\infty$ and $\sup W=+\infty$. Then there exists a minimal complement to $W$.
It remains to deal with the case $\inf W>-\infty$ or $\sup W<+\infty$. Without loss of generality, we may assume that $\inf W>-\infty$ since otherwise we could consider $-W$ instead of $W$. Suppose now that $\inf W=w_{0}>-\infty$, then we can consider the set
$W-(w_{0}-1):=\{w-(w_{0}-1):w\in W\}$
instead of $W$. Thus, without loss of generality, we may further assume that $\inf W=1$. Based on these observations, Chen and Yang [5, Theorem 2] proved the following theorem.
Theorem B (Chen–Yang).
Let $W=\{1=w_{1}<w_{2}<\cdots\}$ be a set of integers and
$\overline{W}=\mathbb{Z}^{+}\setminus W=\{\overline{w_{1}}<\overline{w_{2}}<\cdots\},$
where $\mathbb{Z}^{+}$ is the set of all positive integers as usual.
($a$) Suppose that
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})=+\infty,$
then there exists a minimal complement to $W$.
($b$) Suppose that
$\lim_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty,$
then there does not exist a minimal complement to $W$.
In 2019, Kiss, Sándor and Yang [7] pointed that if
$W=\bigcup\limits_{k=0}^{+\infty}[10^{k},2\times 10^{k}],$
then it clear that
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})=+\infty\quad\text{and}\quad\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty.$
Hence the criterion $\lim_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ in Theorem B($b$) cannot be easily improved to $\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ since otherwise there will be a contradiction between Theorem B($a$) and Theorem B($b$) for this $W$. Therefore, it would be meaningful to think of the following question.
Problem 2.
Let $W=\{1=w_{1}<w_{2}<\cdots\}$ be a set of integers and
$\overline{W}=\mathbb{Z}^{+}\setminus W=\{\overline{w_{1}}<\overline{w_{2}}<\cdots\}.$
Determine the structures of $W$ with $\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})=+\infty$ which does not have a minimal complement.
Extending Theorem B($b$), we answer Problem 2 by the following theorem which can be viewed as a partial solution to Problem 1.
Theorem 1.
Let $W=\{1=w_{1}<w_{2}<\cdots\}$ be a set of integers and
$\overline{W}=\mathbb{Z}^{+}\setminus W=\{\overline{w_{1}}<\overline{w_{2}}<\cdots\}.$
If there exists a sequences $i_{1}<i_{2}<\cdots$ such that
$\lim_{t\rightarrow+\infty}\left(\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}\right)=+\infty\quad\text{and}\quad\limsup_{t\rightarrow+\infty}\big{|}\overline{W}\cap(\overline{w_{i_{t}}},\overline{w_{i_{t+1}}})\big{|}<+\infty,$
then there does not exist a minimal complement to $W$.
It would be of interest to show via the following example that our new theorem has some advantages comparing with Theorem B.
Example 1.
Let
$W=\{1\}\cup\bigcup\limits_{k=4}^{+\infty}\big{[}2^{k}+9,2^{k+1}\big{)}=\{1=w_{1}<w_{2}<\cdots\}.$
So, it is plain that
$\overline{W}=\mathbb{Z}^{+}\setminus W=\bigcup\limits_{k=1}^{+\infty}\big{[}2^{k},2^{k}+8\big{]}=\{\overline{w_{1}}<\overline{w_{2}}<\cdots\}.$
Then, clearly we have
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})\leq 10\quad\text{and}\quad\lim_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})\leq 8.$
Thus, both Theorem B($a$) and Theorem B($b$) fail to determine whether such a $W$ has a minimal additive complement. However, it is easily to see that
$\lim_{k\rightarrow\infty}\Big{(}2^{k+1}-\big{(}2^{k}+8\big{)}\Big{)}=+\infty$
and
$\lim_{k\rightarrow\infty}\big{|}\overline{W}\cap\big{(}2^{k}+8,2^{k+1}+8\big{)}\big{|}=8.$
Therefore, we know that $W$ does not have a minimal additive complement, thanks to the criterion from Theorem 1 with $\overline{w_{i_{t}}}=2^{k}+8$ and $\overline{w_{i_{t+1}}}=2^{k+1}$.
Remark 1.
We now give a simple explanation that the criterion given in Theorem 1 is essentially optimal. We first note that
$\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})\geq\lim_{t\rightarrow+\infty}\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}=+\infty$
for those $W$ assumed in Theorem 1 which are consistent with the condition of Problem 2. We next show that the condition
$\limsup_{t\rightarrow+\infty}\big{|}\overline{W}\cap(\overline{w_{i_{t}}},\overline{w_{i_{t+1}}})\big{|}<+\infty$
cannot be neglected. To see this, let $W=\bigcup\limits_{k=0}^{+\infty}[10^{k},2\times 10^{k}]$. It is clear that for this $W$ there exists a sequences $i_{1}<i_{2}<\cdots$ such that
$\lim_{t\rightarrow+\infty}\left(\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}\right)=+\infty\quad\text{and}\quad\limsup_{t\rightarrow+\infty}\big{|}\overline{W}\cap(\overline{w_{i_{t}}},\overline{w_{i_{t+1}}})\big{|}=+\infty.$
Then there exists a minimal additive complement to $W$ by Theorem B($a$) since $\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})=+\infty$.
Along the same line, Kiss, Sándor and Yang [7] investigated certain sets $W$ with
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})<+\infty$
which have a minimal complement. Precisely, let $X$ be a set of integers. If there exists a positive integer $T$ such that $x+T\in X$ for all sufficiently large integers $x\in X$, then $X$ is said to be eventually periodic with period $T$. If $W$ is eventually periodic with $|\mathbb{N}\setminus W|=+\infty$, then there is some integer $T_{W}$ so that $w+T_{W}\in W$ for sufficiently large $w\in W$, which means that
$\limsup_{i\rightarrow+\infty}(w_{i+1}-w_{i})\leq T_{W}<+\infty.$
So, Kiss, Sándor and Yang focused on eventually periodic sets. Let $m$ be a positive integer and $A$ a set. We appoint that
$A\ ({\rm{mod}}\ m)=\big{\{}i\in[0,m-1]:i\equiv a\ ({\rm{mod}}\ m),a\in A\big{\}}.$
Now, suppose that $W$ is an eventually periodic set with periodic $m$. By shifting a number, we may assume, without loss generality, that $W$ has the following structure:
$W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)},$ (1.1)
where $X_{m}\subseteq\{0,1,\cdots,m-1$}, $Y^{(0)}\subseteq\mathbb{Z}^{-}$, $Y^{(1)}$ are finite sets with $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$ and $\big{(}Y^{(1)}\ ({\rm{mod}}\ m)\big{)}\cap X_{m}=\emptyset$. Here, $\mathbb{Z}^{-}$ represents the set of all negative integers. Kiss, Sándor and Yang [7, Theorems 1 and 2] gave a necessary and sufficient condition, respectively, for eventually periodic sets $W$ which have a minimal additive complement.
Theorem D (Kiss–Sándor–Yang).
Let $W$ be defined in (1.1). If there exists a minimal complement to $W$, then there exists some $C\subseteq\{0,1,\cdots,m-1\}$ such that the following two statements hold:
($a$) $C+\big{(}X_{m}\cup Y^{(1)}\big{)}\ ({\rm{mod}}\ m)=\{0,1,\cdots,m-1\}$;
($b$) For any given $c\in C$, there exists some $y\in Y^{(1)}$ such that
$c+y\not\equiv c^{\prime}+x\ ({\rm{mod}}\ m),$
where $c^{\prime}\in C$ and $x\in X_{m}$.
Theorem E (Kiss–Sándor–Yang).
Let $W$ be defined in (1.1). Suppose that there exists some $C\subseteq\{0,1,\cdots,m-1\}$ such that the following two conditions hold:
($a$) $C+\big{(}X_{m}\cup Y^{(1)}\big{)}\ ({\rm{mod}}\ m)=\{0,1,\cdots,m-1\}$;
($b$) For any given $c\in C$, there exists some $y\in Y^{(1)}$ such that
$c+y\not\equiv c^{\prime}+x\ ({\rm{mod}}\ m),$
where $c^{\prime}\in C\setminus\{c\}$ and $x\in X_{m}\cup Y^{(1)}$.
Then there exists a minimal complement to $W$.
For other related results about minimal additive complements, one may also refer to the works [1, 2, 3, 4, 6, 7, 8, 10, 11, 12].
Our second aim in this article is a new result on the situation
$\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})<+\infty.$
Following (1.1), we write
$W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)},$ (1.2)
where $X_{m}\subseteq\{0,1,\cdots,m-1$}, $Y^{(0)}\subseteq\mathbb{Z}^{-}$ is a finite set with $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$ and $Y^{(1)}$ is a set with $\big{(}Y^{(1)}\ ({\rm{mod}}\ m)\big{)}\cap X_{m}=\emptyset$. So, we would have $\limsup_{i\rightarrow+\infty}(\overline{w_{i+1}}-\overline{w_{i}})<+\infty$ by this setting. It is worth mentioning that in our setting here the set $Y^{(1)}$ does not need to be finite which is slightly more general than the ones considered by Kiss, Sándor and Yang above. Now, we state below our second theorem.
Theorem 2.
Let $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ be as in (1.2). Suppose that $S$ is an integer set satisfying the condition that for any integer set $G$ with property that
$(-\infty,n_{G}]\subseteq G+S$
for some integer $n_{G}$ (probably negative), we have
$G\setminus\{g\}+S=G+S\quad(\forall g\in G).$
Then for any $Y^{(1)}=D+mS$ with $D\subseteq[0,m-1]$ the set $W$ does not have a minimal additive complement.
As an application of Theorem 2, we deduce the following corollary which looks a bit more clean than the theorem itself.
Corollary 1.
Let $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ be as in (1.2) and $k$ a positive integer. If $Y^{(1)}=D+mk\mathbb{N}$ with $D\subseteq[0,m-1]$, then the set $W$ does not have a minimal additive complement.
It is clear that $Y^{(1)}$ in Corollary 1 is not finite. So, it can not be easily decided whether $W$ has a minimal additive complement by Theorem D and Theorem E.
Proof of Corollary 1 via Theorem 2.
Let $G$ be a set satisfying that
$(-\infty,n_{G}]\subseteq G+k\mathbb{N}$
for some integer $n_{G}$. Let $g$ be any fixed element of $G$, we prove that
$n\in G\setminus\{g\}+k\mathbb{N}\quad(\forall g\in G)$
for any $n$, from which it follows that
$(G\setminus\{g\})+k\mathbb{N}=G+k\mathbb{N}.$
In fact, for any $n\in G+k\mathbb{N}$ there exist integers $g^{\prime}\in G$ and $n^{\prime}\in\mathbb{N}$ such that $n=g^{\prime}+kn^{\prime}$. If $g^{\prime}\neq g$, then $n\in G\setminus\{g\}+k\mathbb{N}$. It remains to consider the case $g^{\prime}=g$. In this case, we have
$n=g+kn^{\prime}.$
Clearly, there exists some positive integer $t$ such that $g-kt<n_{G}$. Recall that
$(-\infty,n_{G}]\subseteq G+k\mathbb{N},$
so we have $g-kt\in G+k\mathbb{N}$. Hence, $g-kt=g_{1}+k\ell$ for some positive integer $\ell$ and some $g_{1}\in G$, i.e., $g-k(t+\ell)\in G$. Thus,
$n=g+kn^{\prime}=\big{(}g-k(t+\ell)\big{)}+k(n^{\prime}+t+\ell)\in G\setminus\{g\}+k\mathbb{N}.$
By Theorem 2, there does not exist a minimal complement to $W$.
This completes the proof of Corollary 1. ∎
2. Proof of Theorem 1
Proof of Theorem 1.
Suppose the contrary, we assume that $C$ is a minimal additive complement to $W$. Then we clearly have $\inf C=-\infty$. Let
$C\cap(-\infty,-\overline{w_{i_{1}+1}})=\{c_{1}>c_{2}>\cdots\}.$
For any $s\in\mathbb{Z}^{+}$, we assume that
$\overline{w_{i_{u_{s}-1}+1}}\leq-c_{s}<\overline{w_{i_{u_{s}}+1}}.$ (2.1)
Since $\limsup_{t\rightarrow+\infty}\big{|}\overline{W}\cap(\overline{w_{i_{t}}},\overline{w_{i_{t+1}}})\big{|}<+\infty$, there is an absolute constant $K$ so that
$\limsup_{t\rightarrow+\infty}\big{|}\overline{W}\cap(\overline{w_{i_{t}}},\overline{w_{i_{t+1}}})\big{|}=K,$
which means that there is some $t_{W}$ so that for any $t\geq t_{W}$ we have
$\big{|}\overline{W}\cap(\overline{w_{i_{t}}},\overline{w_{i_{t+1}}})\big{|}\leq K.$ (2.2)
Our proof will be separated into the following three cases.
Case 1. $\liminf_{s\rightarrow+\infty}(-c_{s}-\overline{w_{i_{u_{s}-1}+1}})<+\infty$. Let
$h:=\liminf_{s\rightarrow+\infty}(-c_{s}-\overline{w_{i_{u_{s}-1}+1}}).$
Then $0\leq h<+\infty$. Define
$H:=\{s:-c_{s}-\overline{w_{i_{u_{s}-1}+1}}=h\}.$
It is clear that $H$ is an infinite set since $-c_{s}-\overline{w_{i_{u_{s}-1}+1}}$ are all integers. Let $s_{0}\in H$ be any given integer. Then there exist integers $n_{s_{0}}\in\mathbb{Z}$ and $w_{s_{0}}\in W$ such that
$n_{s_{0}}=c_{s_{0}}+w_{s_{0}}\quad\text{but}\quad n_{s_{0}}\neq c+w$ (2.3)
for any $w\in W$ and $c\in C$ with $c\neq c_{s_{0}}$, otherwise $C\setminus\{c_{s_{0}}\}$ will be an additive complement to $W$ which contradicts with the minimal property of $C$. Recall that $\inf C=-\infty$ and
$\lim_{t\rightarrow+\infty}\left(\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}\right)=+\infty$
from the condition of our theorem. There exists an integer $s_{0}^{*}\geq s_{0}$ such that
$n_{s_{0}}-c_{s}>0$ (2.4)
for all $s\in H$ with $s\geq s_{0}^{*}$ and
$\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}>|n_{s_{0}}|+h$ (2.5)
for all $t\geq u_{s_{0}^{*}}-1$. From (2.3) and (2.4), we have
$n_{s_{0}}-c_{s}\in\overline{W}$ (2.6)
for $s\geq s_{0}^{*}$. From (2.5) with $t=u_{s}$ we have
$\overline{w_{i_{u_{s}}+1}}-\overline{w_{i_{u_{s}}}}>|n_{s_{0}}|+h=|n_{s_{0}}|-c_{s}-\overline{w_{i_{u_{s}-1}+1}}\geq|n_{s_{0}}|-c_{s}-\overline{w_{i_{u_{s}}}}\textsterling\textlnot$
providing that $s\geq s_{0}^{*}$. In other words,
$n_{s_{0}}-c_{s}\leq|n_{s_{0}}|-c_{s}<\overline{w_{i_{u_{s}}+1}}.$ (2.7)
On the other hand, we know again from (2.5) with $t=u_{s}-1$ that
$\overline{w_{i_{u_{s}-1}+1}}-\overline{w_{i_{u_{s}-1}}}>|n_{s_{0}}|+h\geq-n_{s_{0}}.$ (2.8)
So by (2.1) and (2.8), we have
$n_{s_{0}}-c_{s}\geq n_{s_{0}}+\overline{w_{i_{u_{s}-1}+1}}>\overline{w_{i_{u_{s}-1}}},$ (2.9)
provided that $s\geq s_{0}^{*}$. Hence, we conclude from (2.6), (2.7) and (2.9) that
$\overline{w_{i_{u_{s}-1}}}<n_{s_{0}}-c_{s}<\overline{w_{i_{u_{s}}+1}}\quad\text{and}\quad n_{s_{0}}-c_{s}\in\overline{W}$ (2.10)
for all $s_{0},s\in H$ with $s\geq s_{0}^{*}\geq s_{0}$.
Now, we chose $K+2$ different integers $s_{0,1},s_{0,2},...,s_{0,K+2}\in H$. Then there are $K+2$ integers $s_{0,1}^{*},s_{0,2}^{*},...,s_{0,K+2}^{*}$ so that for any $21\leq j\leq K+$ 2 we have
$\overline{w_{i_{u_{s}-1}}}<n_{s_{0,j}}-c_{s}<\overline{w_{i_{u_{s}}+1}}\quad\text{and}\quad n_{s_{0,j}}-c_{s}\in\overline{W}\quad\quad(\forall~{}s\geq s_{0,j}^{*})$ (2.11)
from (2.10). Now let $s$ be a sufficiently large integer so that
$s>\max\left\{s_{0,1}^{*},s_{0,2}^{*},...,s_{0,K+2}^{*}\right\}$
and $u_{s}-1\geq t_{W}$. Then for such $s$ we have
$n_{s_{0,j}}-c_{s}\in\overline{W}\cap(\overline{w_{i_{u_{s}-1}}},\overline{w_{i_{u_{s}}+1}})\quad\quad(\forall~{}1\leq j\leq K+2)$
from (2.11). Moreover, we know from (2.2) that
$\big{|}\overline{W}\cap(\overline{w_{i_{u_{s}-1}}},\overline{w_{i_{u_{s}}+1}})\big{|}=\big{|}\overline{W}\cap(\overline{w_{i_{u_{s}-1}}},\overline{w_{i_{u_{s}}}})\big{|}+\big{|}\overline{W}\cap[\overline{w_{i_{u_{s}}}},\overline{w_{i_{u_{s}}+1}})\big{|}\leq K+1,$
from which it follows that there at least two integers
$21\leq j_{1}<j_{2}\leq K+$ 2
such that
$n_{s_{0,{j_{1}}}}-c_{s}=n_{s_{0,{j_{2}}}}-c_{s},\quad\text{i.e.,}\quad n_{s_{0,{j_{1}}}}=n_{s_{0,{j_{2}}}}.$
By (2.3), we clearly have
$n_{s_{0,{j_{1}}}}=c_{s_{0,{j_{1}}}}+w_{s_{0,{j_{1}}}}\quad\text{and}\quad n_{s_{0,{j_{2}}}}=c_{s_{0,{j_{2}}}}+w_{s_{0,{j_{2}}}},$
which is certainly a contradiction with (2.3) since $n_{s_{0,{j_{1}}}}=n_{s_{0,{j_{2}}}}$ whereas $c_{s_{0,{j_{1}}}}\neq c_{s_{0,{j_{2}}}}$.
Case 2. $\liminf_{s\rightarrow+\infty}(c_{s}+\overline{w_{i_{u_{s}}+1}})<+\infty$. The proof is similar to Case 1. Let
$\ell:=\liminf_{s\rightarrow+\infty}(c_{s}+\overline{w_{i_{u_{s}}+1}}).$
Then $0\leq\ell<+\infty$. Define
$L:=\{s:c_{s}+\overline{w_{i_{u_{s}}+1}}=\ell\}.$
It is clear that $L$ is an infinite set since $c_{s}+\overline{w_{i_{u_{s}}+1}}$ are all integers. Let $s_{0}\in H$ be any given integer. Then there exist integers $n_{s_{0}}\in\mathbb{Z}$ and $w_{s_{0}}\in W$ such that
$n_{s_{0}}=c_{s_{0}}+w_{s_{0}}\quad\text{but}\quad n_{s_{0}}\neq c+w$ (2.12)
for any $w\in W$ and $c\in C$ with $c\neq c_{s_{0}}$, otherwise $C\setminus\{c_{s_{0}}\}$ will be an additive complement to $W$ which contradicts with the minimal property of $C$. Since $\inf C=-\infty$ and
$\lim_{t\rightarrow+\infty}\left(\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}\right)=+\infty$
from the condition of our theorem, there exists an integer $s_{0}^{*}\geq s_{0}$ such that
$n_{s_{0}}-c_{s}>0$ (2.13)
for all $s\in L$ with $s\geq s_{0}^{*}$ and
$\overline{w_{i_{t}+1}}-\overline{w_{i_{t}}}>|n_{s_{0}}|+\ell$ (2.14)
for all $t\geq u_{s_{0}^{*}}$. From (2.12) and (2.13), we have
$n_{s_{0}}-c_{s}\in\overline{W}$ (2.15)
for $s\geq s_{0}^{*}$. By (2.14) with $t=u_{s}+1$, we have
$\overline{w_{i_{u_{s}+1}+1}}-\overline{w_{i_{{u_{s}}+1}}}>|n_{s_{0}}|+\ell\geq n_{s_{0}},$ (2.16)
providing that $s\geq s_{0}^{*}$. From (2.1) and (2.16) for $s\geq s_{0}^{*}$ we have
$n_{s_{0}}-c_{s}<n_{s_{0}}+\overline{w_{i_{u_{s}}+1}}\leq n_{s_{0}}+\overline{w_{i_{u_{s}+1}}}<\overline{w_{i_{u_{s}+1}+1}}.$ (2.17)
On the other hand, using (2.16) again we get
$\overline{w_{i_{u_{s}}+1}}-\overline{w_{i_{u_{s}}}}>|n_{s_{0}}|+\ell=|n_{s_{0}}|+c_{s}+\overline{w_{i_{u_{s}}+1}},$
by the definition of $\ell$, from which it follows that
$n_{s_{0}}-c_{s}\geq-|n_{s_{0}}|-c_{s}>\overline{w_{i_{u_{s}}}},$ (2.18)
provided that $s\geq s_{0}^{*}$. Hence, we conclude from (2.15), (2.17) and (2.18) that
$\overline{w_{i_{u_{s}}}}<n_{s_{0}}-c_{s}<\overline{w_{i_{u_{s}+1}+1}}\quad\text{and}\quad n_{s_{0}}-c_{s}\in\overline{W}$ (2.19)
for all $s_{0},s\in L$ with $s\geq s_{0}^{*}\geq s_{0}$.
Now, the remaining arguments are the same to Case 1, so we just omit them.
Case 3. $\liminf_{s\rightarrow+\infty}(-c_{s}-\overline{w_{i_{u_{s}-1}+1}})=\liminf_{s\rightarrow+\infty}(c_{s}+\overline{w_{i_{u_{s}}+1}})=+\infty.$ Since $C$ is a minimal complement to $W$, then for any given integer $s_{0}\geq 1$ there exist integers $n_{s_{0}}\in\mathbb{Z}$ and $w_{s_{0}}\in W$ such that
$n_{s_{0}}=c_{s_{0}}+w_{s_{0}}\quad\text{but}\quad n_{s_{0}}\neq c+w$ (2.20)
for any $w\in W$ and $c\in C$ with $c\neq c_{s_{0}}$. By conditions of Case 3, there exist infinitely many integers $s>s_{0}$ such that
$-c_{s}-\overline{w_{i_{u_{s}-1}+1}}>|n_{s_{0}}|\quad\text{and}\quad c_{s}+\overline{w_{i_{u_{s}}+1}}>|n_{s_{0}}|,$
which implies that
$n_{s_{0}}-c_{s}\leq|n_{s_{0}}|-c_{s}<\overline{w_{i_{u_{s}}+1}}\quad\text{and}\quad n_{s_{0}}-c_{s}\geq-|n_{s_{0}}|-c_{s}>\overline{w_{i_{u_{s}-1}+1}}.$ (2.21)
Hence, we conclude from (2.20) and (2.21) that
$\overline{w_{i_{u_{s}-1}+1}}<n_{s_{0}}-c_{s}<\overline{w_{i_{u_{s}}+1}}\quad\text{and}\quad n_{s_{0}}-c_{s}\in\overline{W}$
for any $s>s_{0}$.
Using similar arguments of Case 1, the validity of the last case clearly follows. ∎
3. Proof of Theorem 2
Proof of Theorem 2.
Suppose the contrary, we assume that $C$ is a minimal additive complement to $W$. We will deduce a contradiction. For any $i\in[0,m-1]$, we define
$C_{i}=\{c\in C:c\equiv i\ ({\rm{mod}}\ m)\}$
and
$J=\big{\{}j:\big{|}C_{j}\cap\mathbb{Z}^{-}\big{|}=+\infty,~{}0\leq j\leq m-1\big{\}}.$
For any integer (if there does exist) $t\in[0,m-1]\setminus J$, we have $|C_{t}\cap\mathbb{Z}^{-}|<+\infty$, and so
$\inf\bigg{(}\bigcup\limits_{t\in[0,m-1]\setminus J}(C_{t}+W)\bigg{)}>-\infty,$
which implies that
$J+W\ ({\rm{mod}}\ m)=[0,m-1].$
We define $\mathcal{C}_{1}=J+X_{m}\ ({\rm{mod}}\ m)$ and $\mathcal{C}_{2}=J+Y^{(1)}\ ({\rm{mod}}\ m)$. Then
$\mathcal{C}_{1}\cup\mathcal{C}_{2}=J+\big{(}X_{m}\cup Y^{(1)}\big{)}\ ({\rm{mod}}\ m)=J+W\ ({\rm{mod}}\ m)=[0,m-1],$
where the last but one equality follows from the fact $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$. The arguments will be separated into the following two cases.
Case 1. $\mathcal{C}_{2}\subseteq\mathcal{C}_{1}$. Let $c_{0}\in C$ be any fixed number. Noting that $\mathcal{C}_{1}\cup\mathcal{C}_{2}=[0,m-1]$, we clearly have $\mathcal{C}_{1}=[0,m-1]$. So, for any integer $n$, there exist $k\in\mathbb{Z}$, $x\in X_{m}$ and $j\in J$ such that $n=j+x+km$. Recall that $|C_{j}\cap\mathbb{Z}^{-}|=+\infty$ for any $j\in J$. There exists a positive integer $t>|k|$ such that $j-tm\in C$. Thus,
$n=(j-tm)+\big{(}x+(k+t)m\big{)}\in C+W.$
If $j-tm\neq c_{0}$, then $n\in C\setminus\{c_{0}\}+W$. If $j-tm=c_{0}$, then there exists a positive integer $t^{\prime}>t$ such that $j-(t+t^{\prime})m\in C$ since $|C_{j}\cap\mathbb{Z}^{-}|=+\infty$ for any $j\in J$. Thus,
$n=(j-(t+t^{\prime})m)+(x+(k+t+t^{\prime})m)\in C\setminus\{c_{0}\}+W,$
which is a contradiction. The proof of Case 1 is complete by the discussions above.
Case 2. $\mathcal{C}_{2}\not\subseteq\mathcal{C}_{1}$. For any $i\in[0,m-1]$ and $H,K\subseteq[0,m-1]$, we define
$C(i;H,K)=\big{\{}(h,k):h+k\equiv i\ ({\rm{mod}}\ m),~{}h\in H,~{}k\in K\big{\}}.$
For any $(h,k)\in C(i;H,K)$ let
$\mathcal{A}_{h,k}=\big{(}(h+m\mathbb{Z})\cap C\big{)}+\big{(}(k+m\mathbb{Z})\cap W\big{)}.$
Let $\mathscr{C}=C\ ({\rm{mod}}\ m)$ and $\mathscr{W}=W\ ({\rm{mod}}\ m)$. Then for any $i\in[0,m-1]$, we claim
$i+m\mathbb{Z}$ $=\bigcup\limits_{(c,w)\in C(i,\mathscr{C},\mathscr{W})}\mathcal{A}_{c,w}.$ (3.1)
In fact, for any $(c,w)\in C(i,\mathscr{C},\mathscr{W})$ it is plain that
$\mathcal{A}_{c,w}=((c+m\mathbb{Z})\cap C\big{)}+\big{(}(w+m\mathbb{Z})\cap W\big{)}\subseteq(c+m\mathbb{Z})+(w+m\mathbb{Z})\subseteq i+m\mathbb{Z}.$
On the other hand, for any $i+mt\in i+m\mathbb{Z}$, we can assume
$i+mt=c+w\quad(c\in C,w\in W)$
since $C+W=\mathbb{Z}$. We now suppose that $c\equiv c_{1}\ ({\rm{mod}}\ m)$ and $w\equiv w_{1}\ ({\rm{mod}}\ m),$ then
$c_{1}+w_{1}\equiv c+w\equiv i\ ({\rm{mod}}\ m).$
Thus, we have $i+mt\in\mathcal{A}_{c_{1},w_{1}}.$ Note that $W=(m\mathbb{N}+X_{m})\cup Y^{(0)}\cup Y^{(1)}$ with $Y^{(0)}$ being a finite set and $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$. We have $\mathscr{W}=X_{m}+D\ ({\rm{mod}}\ m)$. So, we deduce
$\bigcup\limits_{(c,w)\in C(i,\mathscr{C},\mathscr{W})}\mathcal{A}_{c,w}=\bigg{(}\bigcup\limits_{(c,w)\in C(i,\mathscr{C},X_{m})}\mathcal{A}_{c,w}\bigg{)}\cup\bigg{(}\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\mathcal{A}_{c,d}\bigg{)}.$ (3.2)
Trivially, we have $\mathscr{C}=J\cup(\mathscr{C}\setminus J)$. Hence, we conclude from (3.1) and (3.2) that
$i+m\mathbb{Z}=\!\bigg{(}\bigcup\limits_{(c,w)\in C(i,J,X_{m})}\!\mathcal{A}_{c,w}\bigg{)}\cup\bigg{(}\bigcup\limits_{(c,w)\in C(i,\mathscr{C}\setminus J,X_{m})}\!\mathcal{A}_{c,w}\bigg{)}\cup\bigg{(}\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\!\mathcal{A}_{c,d}\bigg{)}.$ (3.3)
For any $d\in D$, it clear that
$(d+m\mathbb{Z})\cap W=d+mS$
since $D\ ({\rm{mod}}\ m)=Y^{(1)}\ ({\rm{mod}}\ m)$, $Y^{(0)}\ ({\rm{mod}}\ m)\subseteq X_{m}$ and $\big{(}Y^{(1)}\ ({\rm{mod}}\ m)\big{)}\cap X_{m}=\emptyset$. For any $(c,d)\in C(i,\mathscr{C},D)$, we have $c+d\equiv i\ ({\rm{mod}}\ m)$. Suppose that $c+d=i+mk_{c,d}$ and $C\cap(c+m\mathbb{Z})=c+mG_{c}$ for $c\in\mathscr{C}$, then
$\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\mathcal{A}_{c,d}$ $=\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\Big{(}\big{(}(c+m\mathbb{Z})\cap C\big{)}+(d+mS)\Big{)}$
$=\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\Big{(}(c+mG_{c})+(d+mS)\Big{)}$
$=\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\Big{(}(c+d)+m(G_{c}+S)\Big{)}$
$=\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\Big{(}i+mk_{c,d}+m(G_{c}+S)\Big{)}$
$=\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\Big{(}i+m\big{(}(k_{c,d}+G_{c})+S\big{)}\Big{)}$
$=i+m\bigg{(}\Big{(}\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}(k_{c,d}+G_{c})\Big{)}+S\bigg{)}.$ (3.4)
Taking $c_{0}\in C$ with $c_{0}\ ({\rm{mod}}\ m)\in J$. Noting that $C$ is a minimal additive complement to $W$, there exist integers $n_{0}\in\mathbb{Z}$ and $w_{0}\in W$ such that
$n_{0}=c_{0}+w_{0}\quad\text{and}\quad n_{0}\neq c+w$ (3.5)
for any $w\in W$ and $c\in C\setminus\{c_{0}\}$. The remaining argument will be separated into the following two subcases.
Subcase 2.1. If $n_{0}\ ({\rm{mod}}\ m)\in\mathcal{C}_{1}\ ({\rm{mod}}\ m)$, then
$n_{0}\ ({\rm{mod}}\ m)\in J+X_{m}\ ({\rm{mod}}\ m)$
by the definition of $\mathcal{C}_{1}$. Hence, there exist $k\in\mathbb{Z}$, $x\in X_{m}$ and $j\in J$ such that
$n_{0}=j+x+km.$
Noting that $|C_{j}\cap\mathbb{Z}^{-}|=+\infty$ for any $j\in J$, there exists an positive integer $t>|k|$ such that $j-tm\in C$. Thus,
$n_{0}=(j-tm)+(x+(k+t)m)\in C+W.$
If $j-tm\neq c_{0}$, then $n_{0}\in C\setminus\{c_{0}\}+W$. If $j-tm=c_{0}$, then by $|C_{j}\cap\mathbb{Z}^{-}|=+\infty$ again, there exists an positive integer $t^{\prime}>t$ such that $j-(t+t^{\prime})m\in C$. Then
$n_{0}=(j-(t+t^{\prime})m)+(x+(k+t+t^{\prime})m)\in C\setminus\{c_{0}\}+W,$
which leads to a contradiction.
Subcase 2.2. If $n_{0}\ ({\rm{mod}}\ m)\not\in\mathcal{C}_{1}\ ({\rm{mod}}\ m)$, then there exists $i$ with $i\in\mathcal{C}_{2}\setminus\mathcal{C}_{1}$ such that $n_{0}\equiv i\ ({\rm{mod}}\ m)$ since $\mathcal{C}_{1}\cup\mathcal{C}_{2}=[0,m-1]$. It now follows from (3.3) that
$i+m\mathbb{Z}=\bigg{(}\bigcup\limits_{(c,w)\in C(i,\mathscr{C}\setminus J,X_{m})}\!\mathcal{A}_{c,w}\bigg{)}\cup\bigg{(}\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}\!\mathcal{A}_{c,d}\bigg{)}.$ (3.6)
Noting that $|C_{j}\cap\mathbb{Z}^{-}|<+\infty$ for any $j\in\mathscr{C}\setminus J$, there exists an integer $a$ such that
$\bigcup\limits_{(c,w)\in C(i,\mathscr{C}\setminus J,X_{m})}\mathcal{A}_{c,w}\subseteq i+m[a,+\infty),$ (3.7)
which implies from (3), (3.6) and (3.7) that
$(-\infty,n_{G}]\subseteq\bigg{(}\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}(k_{c,d}+G_{c})\bigg{)}+S$ (3.8)
for some integer $n_{G}$. Let $\widetilde{G}_{i}=\bigcup\limits_{(c,d)\in C(i,\mathscr{C},D)}(k_{c,d}+G_{c}).$ Then for any $g\in\widetilde{G}_{i}$ we have
$\widetilde{G}_{i}\setminus\{g\}+S=\widetilde{G}_{i}+S$ (3.9)
thanks to (3.8) and the condition of our theorem.
For the integer $c_{0}$ chosen formerly, we assume that
$c_{0}\equiv c^{\prime}\ ({\rm{mod}}\ m).$
Then there exist $c^{\prime}\in\mathscr{C}$ and $g^{\prime}\in G_{c^{\prime}}$ so that $c_{0}=c^{\prime}+mg^{\prime}$ by the definition of $G_{c^{\prime}}$. Recall that $n_{0}=c_{0}+w_{0}$ with $c_{0}\ ({\rm{mod}}\ m)\in J$ and
$n_{0}\ ({\rm{mod}}\ m)\in\mathcal{C}_{2}(=J+Y^{(1)})\setminus\mathcal{C}_{1}\ ({\rm{mod}}\ m).$
So, we can deduce $w_{0}\in Y^{(1)}.$ Hence, there exist an integer $d\in D$ and $s\in S$ with $(c^{\prime},d)\in C(i,\mathscr{C},D)$ such that $w_{0}=d+ms$. Thus, we have
$n_{0}=$ $c_{0}+w_{0}$
$=$ $(c^{\prime}+mg^{\prime})+(d+ms)$
$=$ $i+mk_{c^{\prime},d}+m(g^{\prime}+s)$
$=$ $i+m((k_{c^{\prime\prime},d}+g^{\prime})+s)$
$=$ $i+m(g+s),$ (3.10)
where $g=k_{c^{\prime},d}+g^{\prime}$. Noting that $g\in\widetilde{G}_{i}$, we have $\widetilde{G}_{i}\setminus\{g\}+S=\widetilde{G}_{i}+S$ by (3.9). Hence, there exist integers $g_{1}\in\widetilde{G}_{i}$ with $g_{1}\neq g$ and $s_{1}\in S$ such that $g+s=g_{1}+s_{1}$. Noting also that $g_{1}\in\widetilde{G}_{i}$, there exists integer $c^{\prime}_{1}\in\mathscr{C}$, integer $d_{1}\in D$ and $g^{\prime}_{1}\in G_{c_{1}^{\prime}}$ such that $g_{1}=k_{c^{\prime}_{1},d_{1}}+g^{\prime}_{1}$ by the definition of $\widetilde{G}_{i}$. Therefore, by (3) we have
$n_{0}=$ $i+m(g_{1}+s_{1})$
$=$ $i+m\big{(}(k_{c^{\prime}_{1},d_{1}}+g^{\prime}_{1})+s_{1}\big{)}$
$=$ $i+mk_{c^{\prime}_{1},d_{1}}+m(g^{\prime}_{1}+s_{1})$
$=$ $(c^{\prime}_{1}+mg^{\prime}_{1})+(d_{1}+ms_{1}).$ (3.11)
Recall that $c^{\prime}_{1}+mg^{\prime}_{1}\in C$ and $n_{0}\neq c+w$ for any $w\in W$ and $c\in C\setminus\{c_{0}\}$ by (3.5), it follows from (3) that
$c_{0}=c^{\prime}+mg^{\prime}=c^{\prime}_{1}+mg^{\prime}_{1},$
which implies that $c^{\prime}=c^{\prime}_{1}$ and $g^{\prime}=g^{\prime}_{1}$. Noting further that $g\neq g_{1}$, we deduce
$g-g_{1}=k_{c^{\prime},d}+g^{\prime}-(k_{c^{\prime}_{1},d_{1}}+g^{\prime}_{1})=k_{c^{\prime},d}-k_{c^{\prime}_{1},d_{1}}\neq 0,$
which leads to the fact $d\neq d_{1}$ because $c^{\prime}=c^{\prime}_{1}$. Since
$c^{\prime}+d\equiv c^{\prime}_{1}+d_{1}\equiv i\ ({\rm{mod}}\ m)$
and $c^{\prime}=c^{\prime}_{1}$, it follows that $d\equiv d_{1}\ ({\rm{mod}}\ m)$. On observing $d,d_{1}\in D$, we have $d=d_{1}$, which is certainly a contradiction. ∎
Acknowledgments
The first named author is supported by National Natural Science Foundation of China (Grant No. 12301003), Anhui Provincial Natural Science Foundation (Grant No. 2308085QA02) and University Natural Science Research Project of Anhui Province (Grant No. 2022AH050171).
The second named author is supported by National Natural Science Foundation of China (Grant No. 12201544), Natural Science Foundation of Jiangsu Province, China (Grant No. BK20210784), China Postdoctoral Science Foundation (Grant No. 2022M710121).